Can someone please help me figure out how to answer this question?How do I solve A=P+PRT FOR A,given that P+$300,r=7%, and t=4 years?
Answer by Annabelle1(69)

A=P+PRT
p=300
r=7%=0.07
t=4
A=300+300*0.07*4=384